OpenStudy (anonymous):

Given ΔABC, m∠A = 60°, b = 9 and c = 6. Find m∠B. Round to the nearest whole degree.

OpenStudy (anonymous):

Sin(<b)/b=Sin(a)/a ... 1st find a.. then findm<b= sin-1(b*sin(a)/a)

OpenStudy (anonymous):

make ur conccepts clear about triangles.. and u will see tahts too easy ,man..